OpenStudy (helpjebal):

How would I go about determining the formula and name of this? A hydrate of Na2CO3 has a mass of 4.31 g before heating. After heating, the mass of the anhydrous compound is found to be 3.22 g.

OpenStudy (welshfella):

we can write the formula for the hydrate as Na2CO3x H2O where x is a whole number and is the number of molecules of water in the hydrate.

OpenStudy (welshfella):

The molecular mass od Na2CO3 = 105.99 and Molecular mass of H2O is 18.008

OpenStudy (welshfella):

in the experiment 4.31 g of the hydrate after heating formed 3.22 g of Na2CO3. The difference in mass is due to the loss of H2O. Mass of H2O lost = 4.31 - 3.22 = 1.09 g Now to find the ratio of molecules of Na2CO3 to water we divide 3.22 and 1.09 by their respective molecular masses 3.22 / 105.99 = 0.03038 1.09 / 18.016 = 0.06050 - which is very close to 1 : what? the second number is the value of x